OpenStudy (anonymous):

i need to find a limit in y=arctg(2x+1) can anyone show me how to do it and tell the answer?

hero (hero):

Try graphing it and see what happens

OpenStudy (anonymous):

\[\text{Let $y = \arctan{(2x+1)}$. Clearly, $x = \frac{1}{2}(\tan{y}-1)$. If $x \to \infty$ it means that}\]\[\text{$\frac{1}{2}(\tan y - 1) \to \infty \Rightarrow \tan y \to \infty \Rightarrow y \to \frac{\pi}{2}$, so $\lim_{x \to \infty} \arctan(2x + 1) = \frac{\pi}{2}.$}\]

hero (hero):

How did you do that Krebante? Anytime I try to write in the equation editor, my letters are always jumbled up together.

hero (hero):

Nevermind, lol

OpenStudy (anonymous):

thanks for the answer, Krebante so much! :)